amandaaaaaaaa:

When comparing the f(x) = –x2 + 2x and g(x) = log(2x + 1), on which interval are both functions positive? (–∞, 0) (0, 2) (2, ∞) (∞, ∞)

Angle:

I can help, give me a minute

amandaaaaaaaa:

omg thank youuuu

Angle:

Here is a picture of both of your equations graphed on the same page https://www.desmos.com/calculator/rzdujsvhya

Angle:

1 attachment
amandaaaaaaaa:

got it

Angle:

You can see that f(x) = -x^2 + 2x is positive between 0 and 2

Angle:

You can also see that g(x) = log (2x + 1) is positive from 0 to infinity

Angle:

so when do you think BOTH are positive?

amandaaaaaaaa:

like when they intersect

Angle:

kind of... more like when are they both above the x axis?

amandaaaaaaaa:

yah thats what they intersect

amandaaaaaaaa:

above the x axis

amandaaaaaaaa:

(1.6,0.6)

Angle:

nono we don't care about intersection

amandaaaaaaaa:

oh okay

amandaaaaaaaa:

sorry

Angle:

f(x) = -x^2 + 2x is positive between 0 and 2 g(x) = log (2x + 1) is positive from 0 to infinity so when are they equations both positive?

amandaaaaaaaa:

when they are above 0

Angle:

correct that is (0, 2) for f(x) and (0, infinity) for g(x) so when do the intervals (0, 2) and (0, infinity) overlap?

amandaaaaaaaa:

(1,1) ?

Angle:



Angle:



amandaaaaaaaa:

okay well that answers my question

amandaaaaaaaa:

Can we do one more

amandaaaaaaaa:

to see if I get it

Angle:

ask separate questions on separate question posts

amandaaaaaaaa:

got it

minesweeper:

to solve this without graphing, you could rewrite f(x) as 1-(x-1)², and if 1-(x-1)²>0 then (x-1)²<1 or -1<x-1<1 and 0<x<2. you also have if g(x)>0 then 2x+1>10⁰=1 so 2x>0 and x>0